An engineer has designed a valve that will regulate water pressure on an automobile engine. The valve was tested on 110 engines and the mean pressure was 7.1 lbs/square inch. Assume the variance is known to be 0.64. If the valve was designed to produce a mean pressure of 7 lbs/square inch, is there sufficient evidence at the 0.01 level that the valve does not perform to the specifications? State the null and alternative hypotheses for the above scenario.

H0 =

Ha =
Answer #1

Solution :

Given that,

Population mean = \mu = 7

Sample mean = \bar x = 7.1

Sample standard deviation = s = 0.8

Sample size = n = 110

Level of significance = \alpha = 0.01

This is a two tailed test.

The null and alternative hypothesis is,

Ho: \mu = 7

Ha: \mu \neq 7

The test statistics,

t = ( \bar x - \mu )/ (s/\sqrt{n})

= ( 7.1 - 7 ) / ( 0.8 / \sqrt{} 110 )

= 1.311

Critical value of  the significance level is α = 0.01, and the critical value for a two-tailed test is

t_{c} = 2.622

Since it is observed that |t| =1.311 \leq t_{c} = 2.622 , it is then concluded that the null hypothesis is fail to rejected.

P- Value = 0.1926   

The p-value is p = 0.1926 > 0.01, it is concluded that the null hypothesis is fail to rejected.

Conclusion :

It is concluded that the null hypothesis Ho is fail to rejected. Therefore, there is not enough evidence to claim that the valve was designed to produce a mean pressure of 7 lbs/square inch, at the 0.01 significance level.